You are here: Foswiki>Commissioning Web>Logs>27Aug013_Log (28 Aug 2013, JuanGuerra)Edit Attach

UT 20130826 AO#2 DX.TS4

Observer:JCGuerra(LBTO),ARiccardi(Arcetri),RBruglio(Arcetri),FQuiros(Arcetri)
Telescope Operator: David(LBTO)
SW Support:
AO Support:
Telescope Support: )
Instrument Support:

Summary

  • Summer shutdown.
  • Calibration of the TS4 at DX side using the 4D.
  • Some problems to startup the adsec.

Detail

Start 9:00pm

Some problem to startup the secondary, the process during the set of the flat didn't finished leaving the coils enabled. The software has been power off two time and the third time the system could be set the flat.

Issue with the communication with the hexapod. The process hexaling ddidn't work, the hexapod can not be moved.

also the adsec eng gui is displaying elevation 2.5 degrees and swing arms 0, this is a wrong configuration, we are at 90 and the swing arm is deployed.

Act #279 removed after FF analisys showed spikes in the residuals

Test of FF/noFF set of the shell

HP in the wrong position, poor alignment

current pos= -0.16 0.94 4.2 -169.6 -141.8 0

new pos= -0.185 0.92 4.21 -157.5 -143 0 new Hexapod intmat calibrated

problems with the communication with the hexapod:

err= hexmoveto(zero_hex_pos) is reading a wrong position

hexmoveto is not working

hexmoveby is not working test of set_flat

test command:

print, fsm_load_program(/auto)

print, fsm_set_flat(/no_flat)

print, fsm_save_shape('shape.sav')

tn='xxx'

opd=get_img4d(nme=3)

save, opd, file=b+tn+'/opd.sav'

data are saved in flat/tracknum: shape.save, opd.sav shape is a fsm_save_shape and contains flattened_status Test Tracknum forces phasemap notes Old FF 20130828_042029 -0.07, 0.02 edge missing @SW no FF 20130828_043053 -0.25 0.1 forces on crate #1

no FF, new d0

20130828_043323

RIP: large bump @NE

the best conditions (WFE and forces) are found with FF and with old d0 calibrations.

we re-enabled the FF in the system (ff_matrix.fits in file configuration .txt) and start acquisition of new FF matrix to improve the last one


Act #279 removed after FF analisys

showed spikes in the residuals Test of FF/noFF set of the shell

HP in the wrong position, poor alignment

current pos= -0.16 0.94 4.2 -169.6 -141.8 0

new pos= -0.185 0.92 4.21 -157.5 -143 0 new Hexapod intmat calibrated

problems with the communication with the hexapod:

err= hexmoveto(zero_hex_pos) is reading a wrong position

hexmoveto is not working

hexmoveby is not working test of set_flat

test command:

print, fsm_load_program(/auto)

print, fsm_set_flat(/no_flat)

print, fsm_save_shape('shape.sav')

tn='xxx'

opd=get_img4d(nme=3)

save, opd, file=b+tn+'/opd.sav'

data are saved in flat/tracknum: shape.save, opd.sav shape is a fsm_save_shape and contains flattened_status Test Tracknum forces phasemap notes Old FF 20130828_042029 -0.07, 0.02 edge missing @SW no FF 20130828_043053 -0.25 0.1 forces on crate #1

no FF, new d0

20130828_043323

RIP: large bump @NE

the best conditions (WFE and forces) are found with FF and with old d0 calibrations.

we re-enabled the FF in the system (ff_matrix.fits in file configuration .txt) and start acquisition of new FF matrix to improve the last one New FF matrix acquisition

print, collect_ff_data(/no_flat)

print, reduce_ff_data(tracknum,/do_plot)

then, from /towerdata/meas/adsec_calib/ff_matrix/20130828_044540/ scp *.fits /home/aoacct/releases/FLAO2/calib/adsec/672a/data/

then make install-calib

20130828_044540. low residual New heaxpod intmat acquisitio Act #279 removed after FF analisys

showed spikes in the residuals Test of FF/noFF set of the shell

HP in the wrong position, poor alignment

current pos= -0.16 0.94 4.2 -169.6 -141.8 0

new pos= -0.185 0.92 4.21 -157.5 -143 0 new Hexapod intmat calibrated

problems with the communication with the hexapod:

err= hexmoveto(zero_hex_pos) is reading a wrong position

hexmoveto is not working

hexmoveby is not working test of set_flat

test command:

print, fsm_load_program(/auto)

print, fsm_set_flat(/no_flat)

print, fsm_save_shape('shape.sav')

tn='xxx'

opd=get_img4d(nme=3)

save, opd, file=b+tn+'/opd.sav'

data are saved in flat/tracknum: shape.save, opd.sav shape is a fsm_save_shape and contains flattened_status Test Tracknum forces phasemap notes Old FF 20130828_042029 -0.07, 0.02 edge missing @SW no FF 20130828_043053 -0.25 0.1 forces on crate #1

no FF, new d0

20130828_043323

RIP: large bump @NE

the best conditions (WFE and forces) are found with FF and with old d0 calibrations.

we re-enabled the FF in the system (ff_matrix.fits in file configuration .txt) and start acquisition of new FF matrix to improve the last one New FF matrix acquisition

print, collect_ff_data(/no_flat)

print, reduce_ff_data(tracknum,/do_plot)

then, from /towerdata/meas/adsec_calib/ff_matrix/20130828_044540/ scp *.fits /home/aoacct/releases/FLAO2/calib/adsec/672a/data/

then make install-calib

20130828_044540. low residual New heaxpod intmat acquisition.

File: 20130828_080353

Pos: -0.12, 0.92,4.2,-165.8,-138.4,0

INTMAT.

AdOpt >print,i4d_opt_intmat(0,250) Tracknum modes aver 20130828_080955 0,250 3 20130828_081132 251,500 3 20130828_081328 501,658 3

data to be discarded as some trigger were lost. trigger interface replaced Other test of load shape with/without FF

same commands as before

data are saved in flat/tracknum: shape.save, opd.sav shape is a fsm_save_shape and contains flattened_status Test Tracknum forces phasemap notes With FF 20130828_082132 no FF 20130828_082505 0.26

Data for the optimization of the modal filtering procedure:

print, fsm_set_flat(/no_fl)

print, fsm_save_shape('shape.sav')

print, fsm_set_flat()

print, fsm_save_shape('shape.sav') data are in /towerdata/meas/adsec_calib/flat/ tracknum / shape.sav Test Tracknum no_flat 20130828_082929 flat 20130828_083116 New set of influence functions Tracknum modes 20130828_105456 0-250 20130828_105629 251-500 20130828_105801 501-658

data have to be discarded because of many recontruction errors. see for instances mode close to 170....

Debugging of 4d parameter

template for the errors: mode 190, 192, 195, 200

20130828_113026

we collect modes in the range 180 --200 and compare the result changing 4d configurations

mode #200 is a reference. an image is saved in /towerdata/meas/adsec_calib/report/mode200-rec-error.jpg Tracknum option result 20130828_105456 template bad 20130828_113622 trim=1 bad 20130828_114006 trim=1, Datafill 1 20130828_114007

same dataset, reproduced changing the modulation threshold

from 0.12 to 0.35 (data copied from original location) same result as before

try also to load an old configuration file!! TEST

mode 200 is applied statically on the mirror, with the same amplitude as during the if sampling. the amplitude is recovered from the associated disturbance file

comm=2e-6*adsec.ff_p_svef

print, fsm_apply_cmd(/delta, /passed, -2.83415e-06*comm)

20 images have been captured with the script: comm4d.capture(20, 'mode200'), that is running @28Hz: data in mode200/hdf5

then a burst is collected with the gui: 20 images, taken every 0s (28hz): data in burst_mode200

the images taken with the scripting are ALL affected by reconstr errors

the images taken with the GUI are ok but a single one, showing a single bad area

same results with the scripting even disabling the data compression and the option in 'Preferences'

-- JuanGuerra - 28 Aug 2013
Topic revision: r2 - 28 Aug 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback